The controller of Hartis Corporation estimates the amount of materials handling overhead cost that should be allocated to the company's two products using the data that are given below:

Wall Mirrors Specialty Windows
Total expected units produced 16,500 1,600
Total expected material moves 1,650 1,500
Expected direct labor-hours per unit 9 4


The total materials handling cost for the year is expected to be $16,562.20.


If the materials handling cost is allocated on the basis of direct labor-hours, the total materials handling cost allocated to the wall mirrors is closest to: (Round your intermediate calculations to 2 decimal places.)

$10,714

$9,065

$14,042

$16,335
Answer #1

Allocation Rate = Total Material Handling Cost/Number of direct labor hours

= 16,562.20/(16500*9+1600*4)

= 16,562.20/154900

= $0.10692188

i.e. $0.11

Hence, allocated to Wall Mirrors = Labor Hours*Rate per hour

= 16,500*9*0.11

= $16,335

Hence, the answer is $16,335